OpenStudy (anonymous):

Simplify 3/x + 13/x-11

hero (hero):

This is adding rational expressions @soccerbabe239. Time to execute some of that knowledge I gave you.

OpenStudy (anonymous):

I got all the way to 16x-33/ x^2-11x im just confused if I need to do more or no?

hero (hero):

\(\large\frac{3}{x} + \frac{13}{x-11} \) \(= \large\frac{3(x - 11)}{x(x-11)} + \frac{13(x)}{x(x-11)}\) \(= \large\frac{3x - 33 + 13x}{x(x - 11)}\) \(=\large\frac{16x - 33}{x(x-11)}\)

hero (hero):

Yes, you got it. That's all that needs to be done

OpenStudy (anonymous):

all I needed to do was simplify the denominator and that's it? sweet! haha thank you @Hero !

hero (hero):

Actually, you could leave it in the form you wrote. Either way, it's correct.

OpenStudy (anonymous):

okay(:

hero (hero):

You still here?